OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? 1.9, 40, 42 2.5, 12, 13 4.6, 9, 10 5.13, 35, 37

Parth (parthkohli):

I just answered a question like this

OpenStudy (anonymous):

Lol

Parth (parthkohli):

Anyway, these are in the form a,b,c. The one which will be the lengths of a right triangle would satisfy \(a^2 + b^2 = c^2\) - This is also known as \(\textbf{PYTHAGOREAN THEOREM}\)

Parth (parthkohli):

@Hi! do you get your answer?

OpenStudy (anonymous):

Which one Is It? I just want the answer :D

OpenStudy (anonymous):

Helping and Girl with her math so i just want the answer

OpenStudy (anonymous):

Answer is 2 5^2+12^2 = 25 +144 =169 169 = 13^2

OpenStudy (anonymous):

The second one is a pythagoran triplet if you know pythagoran triplet then in help you a lo t